Old Service Related Notes

NEWS NOTES (08 Dec 2009): A number of changes have been made. These include:


NEWS NOTE (31 Jul 2008): All image data are now being served from a RAID, no more waits for loading tapes. This should reduce typical image request times to 1 to 2 minutes, instead of 15 to 60 minutes.

NEWS NOTE (28 Mar 2005): Release of Virtual Observatory (VO) based Cone Search tools for accessing the NOMAD, USNO-B1, USNO-A2 and ACT catalogues from this site.

NEWS NOTE (23 Nov 2004): The first release of the NOMAD Catalog is now available.

NEWS NOTE (18 Oct 2002): The USNO-B1.0 Catalog is now available.


Service Alert (11 Dec 2012: 4pm MST): Planned service outage. This server will be taken down briefly for maintenance on 13 Dec 2012 at 4pm MST. The outage is expecdted to last no longer than 10 minutes.

Service Alert (27 Aug 2010: 3pm MST): Service is UP. The troubles with the system disk on the backend server have been fixed.

Service Alert (26 Aug 2010: 8pm MST): Service is DOWN. Troubles with the system disk on the backend server. Hopefully the system will be back on-line tomorrow (27 Aug).

Service Alert (20 Jul 2010: 2:55pm MST): Service Restored. The backend server is back on-line. If you notice problems, please let me know.

Service Alert (19 Jul 2010: 8:30pm MST): Planned service outage. The backend server will be off-line 9am-3pm MST (4pm-10pm UT) on 21 July 2010 for system maintenance work.

Service Alert (30 Mar 2010: 10:10am MST): Planned service outage. The backend server will be off-line 7am-9am MST (2pm-4pm UT) on 31 March 2010 for re-location within the computer room.

Service Update (07 May 2009: 12:30am MST): Service has been restored. The RAID failure led to a complete replacement of the server computer with one of the last of its clones (hurrah for 11 year old components found in the back room), and a complete re-installation of the server software from backups. The system is once again functioning, and I am trying to raise the priority of a full system upgrade to a newer OS and newer hardware. Thanks for your patience.

Service Update (05 May 2009: 6:00pm MST): Service is offline. The primary RAID array for the system has failed again. The catalogue service will be offline until tomorrow (but hopefully not longer). Apologies for the interruption. The hardware is old.

Service Update (31 Mar 2009: 5:00pm MST): Service is back on-line. The RAID has been fixed, and brought back up.

Service Update (29 Mar 2009: 7:20pm MST): Service is offline. The primary RAID array for the system has failed. The catalogue service will be offline until Monday 30 Mary 2009 (but hopefully not longer). Apologies for the interruption.

Service Update (16 Mar 2009: 10:30pm MST): Service brought back up. The NOFS web server was moved to a new machine, and new flavor of operating system. The pixel and catalogue server front end was down while the software was ported. The whole exercise took longer than expected, but was finished within about 14 hours (or so) of the time the change over was initiated. My apologies for the interruption.

Service Update (20 Jan 2009: 8:30pm MST): Service URL change (in progress). The new URL for this service is

http://www.usno.navy.mil/USNO/astrometry/optical-IR-prod/icas
At some point in the near future access through the current web site will cease. I do not know the projected duration of the overlap period. The new web site looks different, but retains all of the functions of the current site, and accesses the same server and data sets.

Service Update (12 Dec 2008: 10am MST): Impending service URL change. When I have a complete URL for this service, I will post it here. Hopefully that will happen before the service moves to the new portal.
NOTICE TO OUR USERS

Under direction from the Chief of Naval Operations, the U.S. Naval Observatory's web site is changing from its traditional appearance to a web "portal" which will combine all of the Navy Oceanography Community's products and services.

As of 15 January 2009, the U.S. Naval Observatory's publicly-accessible products and services will continue to be available at http://www.usno.navy.mil, but will be part of a larger portal.

DoD customers are directed to our NIPRNet site at https://oceanography.navy.mil (CAC required).

Questions or comments can be submitted here.

Service Note (06 Dec 2008: 9pm MST): Impending service URL change. Due to changes in the way the Department of Navy is structuring its web sites, in the near future, the web address for this service will probably change. I do not have any more details at the moment. As I learn more, I will post it.

Service Note (31 Jul 2008: 930am MST): Image service has been restored. The image files for the Southern Surveys (SRC-J, ESO-R, SERC-EJ and SERC-ER) have been loadad onto disk and are available once again from this service. At this point, all the data that were previously available are again available, and with much shorter turn around times, since the tape robot is no longer being used.

I note that the tape robot and drives have put in a decade of service in support of this project. The image server computer is also 10 years old this year. Few are the pieces of modern computing equipment that work this well for this long.

Service Note (30 Jul 2008: 5:25pm MST): More recovery work done to the image system. The image files for POSS-II (IIIaJ, IIIaF and IV-N plates) have been loaded onto a RAID server and are once again available. Work is continuing on the Southern Surveys. (See previous 3 notes).

Service Note (28 Jul 2008: 5:10pm MST): The image system has been partially restored. The image files for the POSS-I plates (both 103aE and 103aO plates) and the AAO-R plates have been loaded onto a RAID that is accessible by this image server. Hence, images for POSS-I and AAO-R are once again being served out. As an added bonus, since all the images are on disk, the time needed to get data back should drop dramatically. I'm still working on the image data for POSS-II and the Southern survey (SRC-J, ESO-R, SERC-EJ, SERC-ER).

Service Note (23 Jul 2008: 8:55am MST): The system is still having troubles with image data. The tape drives are still mis-behaving. They worked ok for about 12 hours and then failed again. Until I can find a more reliable fix for the drives, or I can get some (or all) of the image data onto a disk server usable by this system, images are off-line. The tape robot and drives are now pushing 10 years old, so this slow decline in function is not a complete surprise. My apologies to all who have found this useful. Hopefully I will have this back up in a reliable manner soon (see note from 22 Jul 2008).

Service Note (22 Jul 2008: 10:30pm MST): The system is once again delivering image data properly. During the past week, we had a problem with the tape drives that are used read the tapes that hold the scanned image data. The noticeable result was getting back black images instead of proper images of the sky. The trouble has been fixed. I am working on moving the entire image collection to hard disk drives, which will avoid this problem in the future, and speed up the image access dramatically, but progress is slow.

Service Note (19 Oct 2007 1:10pm MST): THE SYSTEM IS ON-LINE. The RAID disk has been replaced, the server software reinstalled, and the system appears to be working once again. The USNO-A2 catalogue is still in the process of being copied onto the local server disk, but otherwise, catalogue access is functioning. I expect the catalogue copy to finish within the next hour or so.

Service Note (18 Oct 2007 3:00pm MST): THE SYSTEM IS OFF-LINE. Sometime within the last 18 hours, the primary RAIDed disk set that supports this service failed. The old disks have been retired, and a new system is being built now. Sorry for the inconvenience. I hope to have the service back up sometime tonight, or early tomorrow morning.

Service Note (21 Apr 2007 9:00pm MST): BACK ON-LINE! The system is back on-line. I have re-built the server and its ancillary input files. Initial tests show all capabilities are once again functional -- the archive is serving up catalogues, finder charts and image data. At this point, I have not thoroughly tested the system, so I would appreciate hearing about any difficulties, or signs that not everything is functioning correctly.

Service Note (18 Apr 2007 5:50pm MST): The system is still down. Recovery efforts are underway. Expected time of service restoration has been pushed back until Friday 20Apr2007.

Service Note (17 Apr 2007 5:00pm MST): Sometime this morning, the primary RAIDED disk set that supports this service had a serious problem with one of its channels. At this point the disk pack is not accessible, and the service is down. I will try to rebuild it tomorrow (18Apr2007). My apologies for the lack of availability, my time to support this has been quite limited lately. Hopefully the service will be operational by mid-afternoon of 18Apr2007.

Service Note (14-16 Mar 2007 MST): We are expecting to upgrade some of our network hardware over the course of the next few days. There will likely be some interruptions of this service as a result. We will try to keep the interference to a minimum.

Service Note (20 Nov 2006 2:35pm MST): We are expecting to upgrade some of our network hardware over the course of this week (and maybe into next week). There will likely be some interruptions of this service as a result. We will try to keep the interference to a minimum.

Service Note (07 Aug 2006 10:30am MST): As noted in the previous service note, we are undertaking a series of hardware and software upgrades. Today we are in the process of replacing the machine that is our site Web server. I have tested this service on the new machine and operating system, and it all appears to work, BUT there is a good chance someone will notice something that I missed. If so, I would very much appreciate hearing about any troubles you have with this service. The old Web server was running on a machine that was more than 7 years old. It is my sincere hope that the upgrade to more modern technology will make this service faster and more reliable.

Service Note (31 July 2006 10:30pm MST): We are undertaking a variety of hardware and software upgrades over the next several weeks. These may have an impact upon this service. Those of you who have noticed the lack of USNO-B1 catalogue data over the past day or two have seen a symptom of this process. I apologize for the inconvenience, but most of these upgrades are due to site-wide upgrades going on right now. I will do my best to minimize the impact on this service. Hopefully, once these changes are done, things will stabilize, and the service will be a bit more robust thanks to the work on everything around it. This work should also facilitate specific upgrades to this service, a few of which are planned for later this summer, and early this fall.

Service Note (20 Feb 2006 5pm MST): The failed temporary data storage disk on the Web/Ftp server machine has been replaced. The new disk appears to working properly.

Service Note (19 Feb 2006 9:45pm MST):We are having troubles with the temporary data storage disk on this Web/Ftp server machine. The catalogue and image service IS running, but will be down tomorrow (20 Feb 2006) for several hours for repairs to this machine. For those who had troubles earlier today and yesterday, the disk problem is the cause.

On a related note, for those trying to use the USNOFS ftp service, all the data available by FTP (under pub/outgoing) was on the disk that failed. Hence, that service is out until at least tomorrow.

Service Note (31 Jan 2006 6pm MST):Service has been restored. If you have troubles please let us know.

Service Note (31 Jan 2006 3pm MST):Service is still intermittent, work is being done our network. With luck it will be functioning again within a few hours. Sorry for the inconvenience.

Service Note (18 Jan 2006 10:50am MST):Service will be interrupted today for USNOFS hardware maintenance and updates. I can't be sure how long the service will be off-line. With luck it will be functioning again within a few hours. Sorry for the inconvenience.

Service Note (19 Dec 2005 2:00pm MST):The NOMAD catalogue is once again available.

Service Note (18 Dec 2005 1:00am MST):The NOMAD catalogue is temporarily inaccessible due to an internal hardware failure. I hope to have the trouble cleared up by mid-morning MST of 19 Dec 2005. USNO-B1 is the suggested alternative until NOMAD is back on-line. My apologies for any inconvenience this may cause.

Service Note (19 Jan 2005 3:00pm MST):The server and our Web site are back on-line.

Service Note (19 Jan 2005 11:30am MST):The server and our Web site will be off-line for about 1 hour while we change out the batteries in the UPS supporting the server machines. The interruption will occur sometime between 12 noon 2pm MST.

Service Note (02 Dec 2004 5:20pm): Upgrades to the server continue, though things have largely stabilized for now.

Service Note (20 Oct 2004 12noon): The server is in the process being upgraded. With luck, you will not notice anything different. HOWEVER, if you see any strange behavior over the next few days, I would very much appreciate hearing about it. I have tested the modifications on my test server, and hopefully I haven't missed too many bugs, but please be aware. Thank you, and thanks for your patience.

Among the changes you may see, I have modified the catalogue ASCII output format slightly to make it easier to use with some of the new NVO tools, and to ease filtering out the header.

Service Note (14 Jul 2004 - Afternoon): The tape robot is back on-line. Let's hope it keeps working.

Service Note (14 Jul 2004 - Morning): We are having trouble with the tape robot that holds the pixel archive. If you have recently requested images and received images that were all black, then you have been caught by this trouble. I am in the process of trying to fix this. More news as it happens.

Service Note (30 Apr 2004): Minor changes have been made to the catalogue service. The most notable ones are (1) the inclusion of zone-record based id numbers for USNO-A2 catalogue objects and (2) all Flag values are now given as Hex numbers (eg 0x1) and (3) a modest tidying up of the search specification frame. Hopefully none of these changes has introduced any new problems.

Service Note (30 Apr 2004): The service is back on-line.

Service Note (29 Apr 2004): The service is offline for the next 6 to 12 hours. We had a disk failure and the RAID used as cache is being reconstructed now. Sorry for the inconvenience.

Service Note (17 Oct 2003): The service is back on-line.

Service Note (17 Oct 2003): The service is offline for the next 12 to 24 hours. We had a disk failure and the RAID used as cache is being reconstructed now. Sorry for the inconvenience.

Service Note (15 Aug 2003): Power at NOFS was interrupted for most of 15 August 2003. The trouble was caused by several lightning strikes last night along the power lines that come into the site. As of a little before 5pm MST, we are back up and running.

Service Note (09 Jul 2003): Bug in application of precession and/or proper motion corrected. This applied only to the case when a user requested data for a given equinox, with a DIFFERENT epoch.

Service Note (02 Jul 2003): The front end has been upgraded to permit entry of Equinox and Epoch separately (e.g., you can now request equinox J2000, epoch 2004.5). In addition, the precession now applies to the extracted image centers as well (meaning that the images should agree with the catalogues when you request an arbitrary equinox/epoch pair).

Service Note (25 Mar 2003): The front end has been upgraded to provide additional services related to the extraction and selection of objects in the USNO-B1.0 catalogue.

Service Note (20 FEB 2003): THE SERVICE IS UP AGAIN.

Service Note (20 Feb 2003): The service is OFF-LINE until further notice. We have had a multi-disk failure with our RAID system and until that is fixed, the server cannot access the catalogues or its cache space. THIS HAS BEEN FIXED.

Service Note (04 Nov 2002): The web server where returned data are stored (ie, this machine) ran out of disk space on its temporary disk sometime over the weekend of 02-03 November. If you asked for data and received an e-mail confirmation, but were unable to get your data (you got a ``File Not Found'' error message) then most likely your data were caught by this lack of disk space. This has been fixed, and I will try to keep a closer watch on available disk space.

Service Note (23 Oct 2002): Modifications to the USNO-B1.0 service are on-going, but we hope to keep the server up continuously. Some basic documentation on the content of USNO-B has been added (and the catalogue format descriptions for USNO-A and ACT have been updated).

Service Note (18 Oct 2002): As of 18 Oct. 2002, we are serving USNO-B1.0 in addition to all previously served data products. While it may not be immediately apparent, many things have changed behind the scenes to make this possible. If you notice something that appears to no longer work, or looks wrong, I would like very much to hear about it. The help files have NOT YET been updated to reflect the changes. This will happen shortly.

Service Note (25 Sept 2002): The service was unavailable for the period 21 to early 23 September due to changes in our connection to the network. We believe this to be fixed.

Service Note (02 May 2002): We had some trouble with our tape robot late on 30 April 2002. Image data requests between late 30 April, and sometime 01 May may not contain any image data because the robot was not mounting tapes. As of 07:20am MST 02 May 2002, this problem is known to be fixed.

Service Note (30 April 2002): The plate logs for the Southern Surveys (AAO-R and SOUTH = ESO-R, SRC-J, SERC-ER and SERC-EJ) have been updated to correct an error in the listed J2000 plate centers.


Back to Software page
Comments, questions, etc., write to S. Levine
Last updated: 07aug2006 by sel